the angle of elevation to the top of a 10 story skyscraper is measured to be 3 degrees from a point on the ground 2,000 feet from the building. what is the height of the skyscraper to the nearest hundreth foot.


1
Expert's answer
2022-01-27T15:20:00-0500

A representation of the above question is shown below:


"\begin{aligned}
&\text { Angle of elevation, } \theta=3 \degree \\
&\text { Distance between building and observing point }=2000 \text { feet } \\
&\Rightarrow \mathrm{BC}=2000 \text { feet } \\
&\text { To find : Height of the building, } \mathrm{AB} \\
&\text { Since, the building and the ground surface are perpendicular to each other } \\
&\Rightarrow \triangle \mathrm{ABC} \text { is a right angle triangle right angled at } \mathrm{B} \\
&\text { Now, to find height of building using property of tan in } \triangle \mathrm{ABC} \\
&\tan \theta =\frac{\text { Perpendicular }}{B a s e} \\
&\tan 3 \degree=\frac{A B}{2000} \\
&\Longrightarrow 0.0524=\frac{A B}{2000} \\
&\Longrightarrow A B=2000 \times 0.0524 \\
&\Longrightarrow \mathrm{AB}=104.8 \text { feet } \\
&\text { Hence, The height of the building }=104.8 \text { feet } \\
\end{aligned}"